# A 3.3904-Competitive Online Algorithm for List Update with Uniform Costs

### v1

@@ Appendix B Computational Results
This section is based on results of a computational study of the $\textsf{LUP}_{1}$ problem. Our source code is available at [13].
@@
Our computational results serve as sources of insight into the $\textsf{LUP}_{1}$ problem, and our code produces graphs which can be checked for correctness, although the size of the graph may be an obstacle. However, it is important to note that we have not verified the correctness of our results by hand nor do we currently include a computer-assisted formal verification of our results.
@@
B.1 Upper bound of 3 for list length up to 6
We start by discussing our results of a exhaustive computer search for the optimal competitive ratio for the $\textsf{LUP}_{1}$ model for small list lengths:
Computational result.
@@ B.1 Upper bound of 3 for list length up to 6
Our computational method starts with a fixed competitive ratio $3$ and creates a bipartite graph. In one part (we call it the OPT part), we store all pairs $(\sigma,W)$, where $\sigma$ is the current list of the algorithm and $W$ is a work function for this state. Each vertex in this part has $n$ edges, corresponding to the new request $r$. Each edge leads to a vertex in the ALG part of the graph, which stores a triple ($\sigma$, $W^{\prime}$, $r$), where $W^{\prime}$ is the new work function after including the new request $r$ into the sequence. In turn, each ALG vertex has $n!$ edges, each leading to a different permutation $\sigma^{\prime}$, which we can interpret as the new list of the algorithm after serving the request $r$.
@@
B.2 Lower bound for the work function algorithm class
@@
We can translate the algorithms for $\textsf{LUP}_{0}$ from [9] to the $\textsf{LUP}_{1}$ model. For input $\sigma$ to $\textsf{LUP}_{1}$ and a list $\pi$, the work function $W^{\sigma}(\pi)$ defines the optimal cost of serving the input $\sigma$ and ending with the list $\pi$. Then, a work function algorithm, given its current state of the list $\mu$ and a new request $r$, it first serves the request and then reorders the list to end in a position which is a minimizer of $\arg\min_{\pi}W^{\sigma,r}(\pi)+d(\pi,\mu)$, where $d(\pi,\mu)$ is the number of pairwise swaps needed to switch from $\mu$ to $\pi$. The minimizer state is not necessarily unique, hence we talk about the work function algorithm class.
@@
Recall that [9] show that some algorithms from the work function algorithm class are $2$-competitive for $\textsf{LUP}_{0}$, which means they are optimal for that setting. In contrast, we show the following:
Computational result.
@@
The competitive ratio of all algorithms in the work function algorithm class in the partial uniform cost model for $\textsf{LUP}_{1}$ is at least $3.1$.
@@
To back up our claim, we present a lower bound instance graph for list length $5$ where all choices of the work function algorithm class are being considered. This graph is equivalent to the graph construction of Section B.1, with the outedges of the ALG vertices restricted only to all valid choices of the work function algorithm class.
@@
However, since the result of an iterative algorithm is that the potential does not stabilize, we provide a heuristical, tailored method how to produce a lower bound graph from the unstable potentials. A result of this code is then a standard lower bound adversary/algorithm graph, with the adversary presenting a single item in each of its vertices and the competitive ratio of any cycle in the graph being at least $3.1$. The source code to generate this graph can be found in our repository [13].
@@
We remark that list length $3$ and $4$ seem to yield a $3$-competitive algorithm for $\textsf{LUP}_{1}$ even from the work function algorithm class.
Appendix C Lower Bounds on MTF and its Modifications
In this section we show the lower bounds greater than $3$ for several natural algorithms, most of which have previously been considered for this problem. All these bounds (unless explicitly stated) hold for the full cost model, and thus also for the partial one.
Deterministic Bit.
Algorithm Deterministic Bit (Dbit): start with all items unmarked. On request to item $r$, if it is marked, move $r$ to front and unmark it. Otherwise, mark $r$. We show that the competitive ratio of Dbit is at least $3.302$.
To this end, consider a list of length $n$ and denote its items as $x_{0},x_{1},\ldots,x_{n-1}$, counting from the front of the list. We divide the list into two parts, call them $A,B$, where part $A$ has length $cn$ and part $B$ has length $(1-c)n$. Value of the parameter $c$ will be determined later. Formally $A=(x_{0},x_{1},x_{2},\ldots,x_{cn-1})$ and $B=(x_{cn},x_{cn+1},\ldots,x_{n-2},x_{n-1})$. To simplify notation, we denote such order of the list as $A,B$. We define request sequence
$\sigma=($ $x_{cn-1},x_{cn-2},\ldots,x_{1},x_{0},$
$x_{n-1},x_{n-1},x_{n-2},x_{n-2},x_{n-3},x_{n-3},\ldots,x_{cn},x_{cn})^{2}.$
In this sequence, all items of $A$ are requested first, starting from the one furthest in the list. Then every item of $B$ is requested twice, also starting from the furthest one. It results in items of $B$ being transported to the front, thus after the first half of $\sigma$ the order of the list of Dbit is $B,A$. Note, that order of items within $B$ does not change.
Then the same sequence of requests is repeated. The difference is that now requesting any item of $A$ results in moving that item to the front. Then items of $B$ are again requested twice, resulting in final list of algorithm being $B,A$. Costs for Dbit are described below.
$\textsl{{Dbit}}(\sigma)=$ $\frac{cn(cn+1)}{2}$ (access to $A$)
$+(1-c)n\cdot(3n-1)$ (access to $B$ and movement)
$+cn\cdot(2n-1)$ (access to $A$ and movement)
$+(1-c)n\cdot(3n-1)$ (access to $B$ and movement)
$=$ $n^{2}\left(6-4c+\frac{c^{2}}{2}\right)+n\left(\frac{3c}{2}-2\right)$ (by simple math)
Now we define a better algorithm for this sequence, which is an upper bound on Opt. It also starts with list $A,B$. After receiving the first batch of requests to the items of $A$, it moves all the items of $B$ before the items of $A$. Then it doesn’t move anymore until the end of request sequence $\sigma$. Costs of such algorithm are described below.
$\textsl{{Opt}}(\sigma)\leq$ $\frac{cn(cn+1)}{2}$ (access to $A$)
$+cn\cdot(1-c)n$ (swap $A$ with $B$)
$+2\cdot\frac{1+(1-c)n}{2}\cdot(1-c)n$ (access to $B$)
$+\frac{(1-c)n+1+n}{2}\cdot cn$ (access to $A$)
$+2\cdot\frac{1+(1-c)n}{2}\cdot(1-c)n$ (access to $B$)
$=$ $n^{2}\left(2-2c+c^{2}\right)+n\left(2-c\right)$ (by simple math)
Note, that the final order of the list in both algorithms is the same and equal to $B,A$. Moreover, every item is requested an even number of times, so no item is marked after $\sigma$. Thus, this request sequence can be repeated infinitely many times, with sets $A,B$ redefined after each sequence.
With long enough list, $n$ becomes insignificant compared to $n^{2}$, therefore the ratio is $\frac{\textsl{{Dbit}}(\sigma)}{\textsl{{Opt}}(\sigma)}\geq\frac{6-4c+\frac{c^{2}}{2}}{2-2c+c^{2}}$. Let $c=\frac{5-\sqrt{13}}{3}$, then $\frac{\textsl{{Dbit}}(\sigma)}{\textsl{{Opt}}(\sigma)}\geq\frac{3+\sqrt{13}}{2}\geq 3.302$.
Deterministic Bit in partial cost model.
Competitive ratio of Dbit in the partial cost model is at least $4$. We will recursively construct a sequence that achieves this ratio for every length of the list $n\geq 2$. For list of length $n$, we denote a sequence of requests as $\sigma_{n}$. For reasons that will become apparent, we want to index elements from the end of the list, namely the algorithm’s list at the start of the sequence is $x_{n-1},x_{n-2},\ldots,x_{1},x_{0}$.
First, assume $n=2$. Set
$\sigma_{2}=x_{1}\,x_{0}\,x_{0}\,x_{1}\,x_{0}\,x_{0}.$
The cost $\textsl{{Dbit}}(\sigma_{2})=8$. Cost $2$ can be achieved for that sequence by starting with the list order $x_{1},x_{0}$ and swapping elements only after the first request. Note that such algorithm would have the same order of elements after serving the whole sequence as algorithm Dbit.
Now let $n>2$. Set
$\sigma_{n}\;=\;x_{n-1}\,\sigma_{n-1}\,x_{n-1}\,\sigma^{\prime}_{n-1}.$
where by $\sigma^{\prime}_{n-1}$ above we mean the sequence obtained from $\sigma_{n-1}$ by permuting the items $x_{n-2},...,x_{0}$ according to their position in the list after executing $\sigma_{n-1}$. Specifically, if $x_{\pi(n-2)}x_{\pi(n-3)}...x_{\pi(0)}$ is the list after executing $\sigma_{n-1}$ then obtain $\sigma^{\prime}_{n-1}$ from $\sigma_{n-1}$ by replacing each item $x_{i}$ by $x_{\pi(i)}$. For example, for $n=3$, we have
The following observation can be proven by straightforward induction. {observation} In $\sigma_{n}$ each item is requested at least twice, and is requested an even number of times. This implies that after executing request sequence $\sigma_{n-1}$ all items $x_{n-2},x_{n-3},...,x_{0}$ will not be marked and the list will have $x_{n-1}$ at the end.
Let $\textsl{{Dbit}}(\sigma_{n})$ be the cost of Dbit on entire input sequence $\sigma_{n}$ and entire list and let $\textsl{{Dbit}}(\sigma_{n-1})$ be the cost of Dbit on $\sigma_{n-1}$ and list of size $n-1$, disregarding element $x_{n-1}$ ($\textsl{{Opt}}(\sigma_{n})$ and $\textsl{{Opt}}(\sigma_{n-1})$ are defined analogously). Thus, we now calculate impact of $x_{n-1}$ on the cost of Dbit.
At the start of the sequence, element $x_{n-1}$ is in front of the list, marked because of the first request. According to Appendix C, during $\sigma_{n-1}$ every element other than $x_{n-1}$ is requested at least twice. For the first pair of such requests $x_{n-1}$ contributes $3$ to the cost (one per each request and one for the swap when the requested element is move to the front). Thus, additional cost contributed by element $x_{n-1}$ for $\sigma_{n-1}$ is at least $3(n-1)$. Then element $x_{n-1}$ is requested and moved to the front with cost $2(n-1)$. After that we apply Appendix C to $\sigma^{\prime}_{n-1}$ so $x_{n-1}$ contributes additional $3(n-1)$. Thus
The adversary can serve the first $x_{n-1}$ at cost 00 and then move it immediately to the end, so together with the second request to $x_{n-1}$, it will pay $2(n-1)$ to serve it. So in the final list of the adversary $x_{n-1}$ is at the end, just like in the algorithm’s list. Applying this recursively, we obtain that the adversary’s list after serving is the same as the one for the algorithm. This implies that serving $\sigma_{n-1}$ and $\sigma^{\prime}_{n-1}$ will cost at most $2\textsl{{Opt}}(\sigma_{n-1})$.
$\textsl{{Opt}}(\sigma_{n})\;\leq\;2\cdot\textsl{{Opt}}(\sigma_{n-1})+2\cdot(n-1)$ (5)
Half-Move.
Next, we consider the algorithm Half-Move, which moves the requested item to the halfway point between its position and the front of the list. We show that its competitive ratio is at least $6$.
Consider a list of even length $n$. Denote the items as $x_{0},x_{1},\ldots,x_{n-1}$, according to their positions in the initial list of the algorithm. The request sequence consists of $k$ repetitions of requests
$x_{n-1},x_{n-2},x_{n-3},\ldots,x_{\frac{n}{2}}$
for some $k\geq 1$. At the time of the request, the requested item is at the end of the list and it is moved to the middle of the list. Since in the list of even length there are two middle positions, we will assume that the item is always moved to the one further from the front of the list. If it is moved to the other one, the proof only needs to be slightly modified.The cost of any singular request for the algorithm is $n+\frac{n}{2}-1=\frac{3n}{2}-1$. Moreover, after $\frac{n}{2}$ requests, the list returns to its original order. Therefore the total cost of the sequence is $k\cdot\frac{n}{2}\cdot(\frac{3n}{2}-1)$.
It is possible to change the order of the list to $x_{\frac{n}{2}},x_{\frac{n}{2}+1},\ldots,x_{n-1},x_{0},x_{1},x_{2},\ldots,x_{\frac{n}{2}-1}$ before any requests arrive, at cost less than $n^{2}$. Then, with no further movement, total cost of the request sequence is $k\cdot\binom{\frac{n}{2}}{2}$. Thus, the total optimal cost is upper-bounded by $n^{2}+\frac{kn^{2}}{8}$. The ratio approaches $6$ for sufficiently large $n$, as $k\rightarrow\infty$.
C.1 Stay-or-MTF Class
Here we consider a class of online algorithms that never make partial moves. That is, when an item $r$ is requested, the algorithm either does not change the position of $x$ or moves $x$ to the front of the list. We refer to such algorithms as Stay-or-MTF algorithms.
We show that if ${\cal{A}}$ is a Stay-or-MTF algorithm then the competitive ratio of ${\cal{A}}$ is at least $3.25$, even for a list of length $3$.
The adversary strategy is illustrated in the figure below. The list has three items named $a$, $b$ and $c$. The vertices represent states of the game. In each state we assume that the algorithm’s list is $abc$. When an algorithm executes a move, the items are appropriately renamed. Each state is specified by the current offset function. This offset function is equal to the current work function minus the minimum of this work function. At each move, the increase of the minimum of the work function is charged towards the adversary cost, and the offset function is updated. Each offset function that appears in this game is specified by the set of permutations of $a,b,c$ where its value is 00. (Its value at each other list is equal to the minimum swap distance to some offset function with value 00.) To specify this set, we inducate, using braces, which elements are allowed to be swapped. For example, $a{\left\{{bc}\right\}}$ represents two permutations, $abc$ and $acb$, and ${\left\{{abc}\right\}}$ represents all permutations.
The adversary strategy is represented by the transitions between the states. For each state the adversary specifies the request she makes when this state is reached. Notation $r/\xi$ on an arrow means that the request is $x$ and the adversary cost is $\xi$. Then the algorithm decides to make a move. The decision points of the algorithm are indicated by small circles. The type of move is uniquely specified by the cost, specified on the edges from these circles. For example, if the request is on $c$, the algorithm either stays, which costs $2$, or moves to front, which costs $4$.
The numbers next to each state are the potential values. We claim that for each move we have
$\Delta{\cal{A}}+\Delta\Phi\;\geq\;\textstyle{\frac{13}{4}}{\cdot}\Delta\textsl{{Opt}}\,,$
where $\Delta{\cal{A}}$, $\Delta\Phi$ and $\Delta\textsl{{Opt}}$ represent the cost of ${\cal{A}}$, the potential change, and the optimum cost for this move. This inequality can be verified by routine inspection of the transition graph. This shows that the competitive ratio of ${\cal{A}}$ is at least ${\frac{13}{4}}$.
It can be also shown that for $n=3$ ratio ${\frac{13}{4}}$ can be achieved with a Stay-or-MTF algorithm. For $n=4$ we can improve the lower bound to $3.3$, using an approach that is fundamentally similar to the one above, but requiring many more states.
Appendix D A 3-Competitive Algorithm for Lists of Length 3
The Fpm algorithm is $3$-competitive for lists of length 3. This is true even against our lower bound $\sum_{x,y}w^{xy}$ on the optimum cost.
To prove that, we consider possible states of Fpm. There are 15 of them. We assign a potential function $\Psi$ to every state and show that for any state and for any request it holds that $\Delta\textsl{{Fpm}}+\Delta\Psi\leq 3\times\sum_{\{x,y\}\in\mathcal{P}}\Delta w^{xy}$. Note that $\Psi$ is introduced only to simplify the proof and is independent from the algorithm.
The results are shown in Table 4. In each state we assume that the list is abc, and the items are renamed after each transition, if needed. Arrows represent modes: an arrow pointing from an item $x$ towards an item $y$ indicates that the pair $\{x,y\}$ is in mode $\alpha$ (if $x$ is before $y$ in the list) or $\gamma$ (otherwise). Otherwise the pair is in mode $\beta$. Column “MOVE” specifies what Fpm does after the request. Column “DEST.” denotes destination state after movement. State changes corespond to moves of Fpm. Cost of pair-based Opt is in the last column: $P_{OPT}=\sum_{\{x,y\}\in\mathcal{P}}\Delta w^{xy}$.
STATE NUMBER $\Psi$ REQ. MOVE DEST. $\Delta\textsl{{Fpm}}$ $\Delta\Psi$ $P_{OPT}$
[Uncaptioned image] 00 00 $a$ Front 00 00 00 00
$b$ $\Theta_{b}$ $1$ $1$ ${\frac{1}{2}}$ ${\frac{1}{2}}$
$c$ $\Theta_{c}$ $2$ $2$ $1$ $1$
[Uncaptioned image] $1$ ${\frac{1}{2}}$ $a$ Front 00 00 $-\textstyle{\frac{1}{2}}$ ${\frac{1}{2}}$
$b$ Front 00 $2$ $-\textstyle{\frac{1}{2}}$ ${\frac{1}{2}}$
$c$ $\Theta_{c}$ $3$ $2$ $1$ $1$
[Uncaptioned image] $2$ $1$ $a$ Front $4$ 00 $-\textstyle{\frac{1}{2}}$ ${\frac{1}{2}}$
$b$ $\Theta_{b}$ $5$ $1$ $1$ $1$
$c$ Front 00 $4$ $-1$ $1$
[Uncaptioned image] $3$ $\frac{3}{2}$ $a$ Front $4$ 00 $-1$ $1$
$b$ Front $4$ $2$ $-1$ $1$
$c$ Front $4$ $4$ $-1$ $1$
[Uncaptioned image] $4$ ${\frac{1}{2}}$ $a$ Front $4$ 00 00 00
$b$ $\Theta_{b}$ $1$ $1$ 00 $1$
$c$ $\Theta_{c}$ $1$ $3$ 00 $1$
[Uncaptioned image] $5$ $2$ $a$ Front $6$ 00 00 $1$
$b$ Front $4$ $2$ $-\frac{3}{2}$ ${\frac{1}{2}}$
$c$ Front $7$ $4$ $-1$ $1$
[Uncaptioned image] $6$ $2$ $a$ Front $6$ 00 00 00
$b$ $\Theta_{b}$ $1$ $1$ $-\frac{3}{2}$ ${\frac{1}{2}}$
$c$ $\Theta_{c}$ $8$ $3$ 00 $1$
[Uncaptioned image] $7$ $1$ $a$ Front $4$ 00 $-\textstyle{\frac{1}{2}}$ ${\frac{1}{2}}$
$b$ $\Theta_{b}$ $9$ $1$ $1$ $1$
$c$ Front 00 $4$ $-1$ $1$
[Uncaptioned image] $8$ $2$ $a$ Front $10$ 00 $-\textstyle{\frac{1}{2}}$ ${\frac{1}{2}}$
$b$ Front 00 $2$ $-2$ $1$
$c$ $\Theta_{c}$ $11$ $2$ $1$ $1$
[Uncaptioned image] $9$ $2$ $a$ Front 00 00 $-2$ $1$
$b$ Front $10$ $2$ $-\textstyle{\frac{1}{2}}$ ${\frac{1}{2}}$
$c$ $\Theta_{c}$ $12$ $2$ $1$ $1$
[Uncaptioned image] $10$ $\frac{3}{2}$ $a$ Front $10$ 00 00 00
$b$ $\Theta_{b}$ $1$ $1$ $-1$ $1$
$c$ $\Theta_{c}$ $13$ $2$ $1$ $1$
[Uncaptioned image] $11$ $3$ $a$ Front $14$ 00 $-1$ $1$
$b$ Front $7$ $2$ $-2$ $1$
$c$ Front $4$ $4$ $-\frac{5}{2}$ ${\frac{1}{2}}$
[Uncaptioned image] $12$ $3$ $a$ Front $7$ 00 $-2$ $1$
$b$ Front $14$ $2$ $-1$ $1$
$c$ Front $4$ $4$ $-\frac{5}{2}$ ${\frac{1}{2}}$
[Uncaptioned image] $13$ $\frac{5}{2}$ $a$ Front $14$ 00 $-\textstyle{\frac{1}{2}}$ ${\frac{1}{2}}$
$b$ $\Theta_{b}$ $3$ $1$ $-1$ $1$
$c$ Front 00 $4$ $-\frac{5}{2}$ ${\frac{1}{2}}$
[Uncaptioned image] $14$ $2$ $a$ Front $14$ 00 00 00
$b$ $\Theta_{b}$ $8$ $1$ 00 $1$
$c$ $\Theta_{c}$ $1$ $3$ $-\frac{3}{2}$ ${\frac{1}{2}}$
Table 4: Fpm is 3-competitive for lists of length 3
There exists an easier $3$-competitive algorithm for $3$ items. This algorithm uses $6$ states to keep track of the past history. The algorithm is illustrated below as a graph whose vertices represent the current state of the algorithm and edges represent its transitions. Again, items are renamed after each transition so that the list is always abc. The modes of each pair are represented the same way as in Table 4. Each edge is labeled by the requested item, the algorithm’s cost, and the optimum cost (${\frac{1}{2}}$ for each mode change). The algorithm’s move is implied by the cost value. For example, if the request is c and the cost is $3$, it means that the algorithm paid $2$ to access c and then swapped c with the preceding b.
margin: MBa: Here $\Phi$ is not the potential used in the proof. Maybe rename to $\Psi$ to avoid confusion and be consistent with Table 4?
Appendix E Lower bound for the Full-Or-Partial-Move algorithm
In this section, we show that for the choice of parematers from 4.12, the competitive ratio of Fpm is at least $3.04$ in the partial cost model.
We construct a sequence that achieves that ratio. That sequence, shown in Table 5, uses only five items, but can be used on any longer list, without requests to any items further than the fifth. Graphs in column “Work function on pairs” represent modes of all the pairs. Arrow from item $x$ to $y$ means that $W^{xy}(xy)+1=W^{xy}(yx)$ (mode $\alpha$ or $\gamma$, depending on the order of the list). Some arrows are omitted for succinctness - the true graph is always the transitive closure of the graph shown.
Table 5 compares cost of Fpm to the cost of pair-based Opt defined as $P_{OPT}=\sum_{\{x,y\}\in\mathcal{P}}\Delta w^{xy}$, which for this sequence is equal to $25$. We now show the behaviour of an offline algorithm that achieves this cost. Before the start of the sequence, the algorithm changes its list to $cdeab$. This incurs some additional one time cost, which is negligible if the sequence is repeated sufficiently many times. During the entire sequence, the algorithm performs two moves: after serving the first request, it moves item $c$ to the third position in the list and after ninth request it brings item $c$ back to the front. Thus, the total cost of swaps for the offline algorithm is $4$. It is easy to check that the cost of requests is $21$, so the total cost is $25$.
List of Fpm Work function Req. Move Cost $P_{OPT}$
on pairs of Fpm of Fpm
$a$ $b$ $c$ $d$ $e$ $\theta_{a}$ $\theta_{b}\ \theta_{d}$ $\theta_{c}$ $\theta_{e}$ [Uncaptioned image]
[…]
$a$ $b$ $c$ $d$ $e$ $\theta_{a}\ \theta_{c}$ $\theta_{b}\ \theta_{d}$ $\theta_{e}$ [Uncaptioned image]
[…]
$a\ \ \theta_{a}$ $b$ $c$ $d$ $e$ $\theta_{c}\ \theta_{e}$ $\theta_{b}\ \theta_{d}$ [Uncaptioned image]
[…]
$e$ $a$ $b$ $c$ $d$ $\theta_{e}$ $\theta_{a}\ \theta_{c}$ $\theta_{b}\ \theta_{d}$ [Uncaptioned image]
[…]
$e$ $a$ $d$ $b$ $c$ $\theta_{e}\ \theta_{d}$ $\theta_{a}\ \theta_{c}$ $\theta_{b}$ [Uncaptioned image]
[…]
$c$ $e$ $a$ $d$ $b$ $\theta_{c}$ $\theta_{d}\ \theta_{e}$ $\theta_{a}$ $\theta_{b}$ [Uncaptioned image]
$d$ $\theta_{d}$ 33 11
$c$ $d$ $e$ $a$ $b$ $\theta_{c}\ \theta_{d}$ $\theta_{e}$ $\theta_{a}$ $\theta_{b}$ [Uncaptioned image]
$e$ $\theta_{e}$ 35 12
$c\ \ \theta_{c}$ $d$ $e$ $a$ $b$ $\theta_{d}\ \theta_{e}$ $\theta_{a}$ $\theta_{b}$ [Uncaptioned image]
[…]
$e$ $c$ $d$ $a$ $b$ $\theta_{e}$ $\theta_{c}\ \theta_{d}$ $\theta_{a}$ $\theta_{b}$ [Uncaptioned image]
[…]
$e$ $d$ $c$ $a$ $b$ $\theta_{d}\ \theta_{e}$ $\theta_{c}$ $\theta_{a}$ $\theta_{b}$ [Uncaptioned image]
[…]
$e\ \ \theta_{c}$ $d$ $c$ $a$ $b$ $\theta_{d}\ \theta_{e}$ $\theta_{a}$ $\theta_{b}$ [Uncaptioned image]
[…]
$c$ $e$ $d$ $a$ $b$ $\theta_{c}$ $\theta_{d}\ \theta_{e}$ $\theta_{a}$ $\theta_{b}$ [Uncaptioned image]
[…]
$c$ $e$ $d$ $a$ $b$ $\theta_{b}\ \theta_{c}$ $\theta_{d}\ \theta_{e}$ $\theta_{a}$ [Uncaptioned image]
[…]
$b$ $c$ $e$ $d$ $a$ $\theta_{b}$ $\theta_{c}$ $\theta_{d}\ \theta_{e}$ $\theta_{a}$ [Uncaptioned image]
$d$ $\theta_{d}$ 64 21
$b$ $c$ $d$ $e$ $a$ $\theta_{b}\ \theta_{d}$ $\theta_{c}$ $\theta_{e}$ $\theta_{a}$ [Uncaptioned image]
$a$ $\theta_{a}$ 68 23
$b\ \ \theta_{a}$ $c$ $d$ $e$ $a$ $\theta_{b}\ \theta_{d}$ $\theta_{c}$ $\theta_{e}$ [Uncaptioned image]
$a$ Front 76 25
$a$ $b$ $c$ $d$ $e$ $\theta_{a}$ $\theta_{b}\ \theta_{d}$ $\theta_{c}$ $\theta_{e}$ [Uncaptioned image]

### v3

@@ Appendix E Computational Results
This section contains the description of our computational study of the $\textsf{LUP}_{1}$ problem. Our source code for reproducing the results is available at [11].
@@
Our computational results serve as sources of insight into the $\textsf{LUP}_{1}$ problem, and our code produces graphs which can be checked for correctness, although the size of the graph may be an obstacle.
@@
E.1 Upper bound of 3 for list length up to 6
We start by discussing our results of a exhaustive computer search for the optimal competitive ratio for the $\textsf{LUP}_{1}$ model for small list lengths.
Computational Result E.1.
@@ E.1 Upper bound of 3 for list length up to 6
Our computational method starts with a fixed competitive ratio $3$ and creates a bipartite graph. In one part (we call it the Opt part), we store all pairs $(\sigma,W)$, where $\sigma$ is the current list of the algorithm and $W$ is a work function for this state. (Recall that the work function $W$ is the optimal cost of serving the input $\sigma$ and ending in a given list state.) Each vertex in this part has $n$ edges, corresponding to the new request $r$. Each edge leads to a vertex in the Alg part of the graph, which stores a triple ($\sigma$, $W^{\prime}$, $r$), where $W^{\prime}$ is the new work function after including the new request $r$ into the sequence. In turn, each Alg vertex has $n!$ edges, each leading to a different permutation $\sigma^{\prime}$, which we can interpret as the new list of the algorithm after serving the request $r$.
@@
E.2 Lower bound for the work function algorithm class
@@
We can translate the algorithms for $\textsf{LUP}_{\mathrm{S}}$ from [9] to the $\textsf{LUP}_{1}$ model. For input $\sigma$ to $\textsf{LUP}_{1}$ and a list $\pi$, the work function $W^{\sigma}(\pi)$ defines the optimal cost of serving the input $\sigma$ and ending with the list $\pi$. Then, a work function algorithm, given its current state of the list $\mu$ and a new request $r$, it first serves the request and then reorders the list to end in a state $\pi$ being a minimizer of $W^{\sigma,r}(\pi)+d(\pi,\mu)$, where $d(\pi,\mu)$ is the number of pairwise swaps needed to switch from $\mu$ to $\pi$. The minimizer state is not necessarily unique, hence we talk about the work function algorithm class.
@@
Recall that some algorithms from the work function algorithm class are $2$-competitive for $\textsf{LUP}_{\mathrm{S}}$ [9], which means they are optimal for that setting. In contrast, we show the following.
Computational Result E.2.
@@
The competitive ratio of all algorithms in the work function algorithm class for $\textsf{LUP}_{1}$ is at least $3.1$ (for the partial cost model).
@@
To show this claim, we construct a lower bound instance graph for list length $5$ where all choices of the work function algorithm class are being considered. This graph is equivalent to the graph construction of subsection E.1, with the out-edges of the Alg vertices restricted only to all valid choices of the work function algorithm class.
@@
However, since the result of an iterative algorithm is that the potential does not stabilize, we constructed a tailored approach for producing a lower bound graph from the unstable potentials. A result of this code is then a standard lower bound using an adversary/algorithm graph, with the adversary presenting a single item in each of its vertices and the competitive ratio of any cycle in the graph being at least $3.1$.
@@
We remark that this approach yields a $3$-competitive work function algorithm for $\textsf{LUP}_{1}$ for list lengths $3$ and $4$.